OpenStudy (anonymous):

factor by grouping.assume any variable exponents represent whole numbers. x^3+ax^2-3x-12

OpenStudy (aum):

Is the coefficient of x^2 'a' ? x^3 + ax^2 - 3x - 12 = (x^3 + ax^2) - (3x + 12) = x^2(x + a) - 3(x + 4) I am not sure if this is what they want.